Martin has $200 in his savings account.He wants to save more money. He is looking at two investment plans. Underplan A, he will increase his accountbalance by $40 a year. Under plan B, hewill increase his account balance by 15%each year. How much more will he save with Plan B after 10 years?
Answer by jorel555(1290)

Under plan A, Martin would have:
200+40(10)=600 dollars in his account.
Plan B:
200 x (115%)^10=809.11
809.11-600= 209.11 more dollars in his account under plan B!!!!!!!